# Measuring Diversity in Synthetic Datasets

### v1

@@ D.2.1 Hyperparameter Settings of Diversity Evaluation
  • DCScore. We employ the cosine similarity as $\operatorname{Kernel}(\cdot)$, and Softmax as $f_{\textbf{K}}(\cdot)$. Except for hyperparameter sensitivity experiments, we set $\tau$ in (4) to 1 for all other experiments.
@@
  • VendiScore. We employ the cosine similarity as $\operatorname{Kernel}(\cdot)$.
D.2.2 Hyperparameter Settings of Downstream Task Training
@@ D.2.3 Evaluation protocol
D.2.4 Prompt Settings of LLM Evaluation
In Section 5, we use GPT-4 to perform pairwise diversity comparisons. To guide GPT-4 in making these comparisons, we employ a well-designed prompt, as illustrated in Figure 6. The prompt for GPT-4 evaluations includes the task definition, diversity definition, general prompt, and sentence sets to be compared.
@@ D.2.4 Prompt Settings of LLM Evaluation
Appendix E Additional Experiments
E.1 Computational Cost for larger datasets
In the case where the inner product is used as the kernel function and $n\gg d$, VendiScore can significantly reduce computational complexity. To ensure a fair comparison, we compare the computation times of VendiScore and DCScore on larger datasets, as well as under different kernel functions. Specifically, we employ SST2, Yelp, and AG News as the evaluated datasets. We randomly sample 4k, 8k, 16k, 32k, 64k samples and record the computation times for both methods across these different sample sizes. We repeat the experiments 5 times to report the mean and standard deviation.
As shown in Tables 69, and 10, DCScore has a shorter computation time than VendiScore in most cases, with VendiScore only exhibiting a computational advantage when the inner product is used and $n\gg d$. Furthermore, as the sample size increases, the efficiency advantage of DCScore becomes more pronounced. When using a polynomial kernel, on the SST2 dataset, DCScore requires only one-third of the computation time of VendiScore when the sample size reaches 64k. In contrast, although VendiScore has a computational advantage in the case of the inner product, the difference compared to DCScore is not significant. The experimental results are consistent with our complexity analysis presented in Section 4.3. Overall, DCScore outperforms VendiScore in terms of computation time across most kernel functions. VendiScore exhibits a computational time advantage only when the inner product is used as the kernel function, which will limit its applicability. As shown in Chapter 4 of (Seeger, 2004), it is essential to employ different kernel functions to accommodate a wider range of scenarios.
@@ E.1 Computational Cost for larger datasets
E.2 Correlation with Downstream Task Training
To investigate the correlation between DCScore and downstream task training, we train text classification models using self-generated datasets under zero-shot and few-shot settings. We vary the generation temperature $\tau_{g}$ of self-generated datasets within the range of $\{0.2,0.7,1.2\}$. More details of training datasets and hyperparameters are presented in Appendix D.1.1 and D.2.2, respectively. Figure 7 shows the loss curves of these trained classification models. In the zero-shot setting, we observe increasing optimal loss values as $\tau_{g}$ varied from 0.2 to 1.2, indicating that the model is more easily fitted to datasets with limited diversity. However, as shown in Table 7, models trained on more diverse datasets achieve better accuracy, which can be attributed to their enhanced generalization capabilities. From Table 7, the diversity evaluated by DCScore has a similar trend to the accuracy performance in the zero-shot setting, further demonstrating the effectiveness of DCScore in diversity evaluation.
@@ E.2 Correlation with Downstream Task Training
In the few-shot setting, we observe a trend in optimal loss variation similar to that in the zero-shot setting, as shown in Figure 7. However, as shown in Table 7, the performance of the model trained on the dataset generated with $\tau_{g}=1.2$ is inferior to that of the model trained on the dataset generated with $\tau_{g}=0.7$, which contrasts with the findings in the zero-shot setting. This phenomenon can be attributed to the higher diversity of datasets generated at a higher $\tau_{g}$, resulting in increased fitting difficulty. Under the current settings, the number of training epochs for the dataset generated at a temperature of 1.2 is insufficient, preventing the trained model from achieving optimal performance. To validate this hypothesis, we increase the number of epochs to 240 and 360 and train models on the dataset generated at a temperature of 1.2. The final training loss and accuracy of these models are shown in Figure 8. We observe that as the number of epochs increases, the model’s loss gradually decreases, and its performance improves progressively. Ultimately, the model’s accuracy outperforms that of models trained on datasets generated at temperatures of 0.2 and 0.7. Moreover, from Table 7, models trained on datasets from the zero-shot setting outperform those trained on datasets from the few-shot setting. However, this discrepancy arises from the different test sets used in the two settings, making direct performance comparisons inappropriate.
@@
E.3 Impact of Embedding Functions $\Phi$
The paradigm of the transformation-based method enables DCScore to utilize various embedding functions tailored to different scenarios. Consequently, we investigate the impact of embedding functions on self-generated datasets used in Section 5.2.1. As shown in Table 11, we compare the correlation of the diversity evaluation results of DCScore across 4 different embedding functions with diversity pseudo-truths, where the model names in parentheses within the embedding function refer to those available on Hugging Face. Our findings indicate that DCScore exhibits strong correlations with diversity pseudo-truths across various embedding functions. Notably, DCScore utilizing the BGE embedding function achieves the best results in half of the cases. Additionally, the minimum correlation in Table 11 exceeds 0.96, which is classified as a strong correlation according to (Akoglu, 2018). This result also supports the following two conclusions: (1) the embedding function used effectively captures the differences among samples in the dataset from multiple perspectives, and (2) DCScore is sufficiently adaptable to different embedding functions while maintaining stable performance.
Table 11: Correlation (Spearman’s $\rho$) results of DCScore with various embedding models. Spearman’s $\rho$ varies between -1 and +1 with 0 implying no correlation. Best results are indicated in bold.
@@ E.3 Impact of Embedding Functions $\Phi$
E.4 Impact of Kernel Functions
Similar to Appendix E.3, we investigate the impact of different kernel functions on the performance of DCScore. Specifically, this experimental setup is identical to that in Appendix E.3. As shown in Table 12, we find that DCScore demonstrates stable performance across various kernel functions. However, the influence of the kernel function is slightly more pronounced than that of the embedding function, as indicated by the greater fluctuations in correlation among the different kernel functions. Furthermore, we observe that DCScore achieves optimal performance in the case of the inner product. Overall, DCScore consistently maintains strong diversity evaluation performance across different kernel functions.
Table 12: Correlation (Spearman’s $\rho$) results of DCScore with various kernel functions. Spearman’s $\rho$ varies between -1 and +1 with 0 implying no correlation. Best results are indicated in bold.
@@ E.4 Impact of Kernel Functions
Appendix F Detailed Modeling of Existing Methods
We present the detailed modeling of existing methods into DCScore as follows:
Distinct-n. Distinct-n (Li et al., 2015) is a prevalent diversity metric depending on n-grams, where n signifies the number of successive items. Distinct-n calculates the proportion of unique n-grams to all n-grams. The n-grams operation falls under the text representation stage, while the step of obtaining a unique set of n-grams corresponds to the pairwise similarity stage. Typically, a high form similarity among samples in the evaluated dataset results in a smaller unique n-grams set. Finally, ratio calculations belong to the diversity summarization stage.
@@ Appendix F Detailed Modeling of Existing Methods
K-means inertia. K-means inertia (Du & Black, 2019), a transformation-based method, performs clustering in sample representation and then calculates inertia as diversity outcomes. Here, inertia refers to the square summarization of the distance between samples and cluster centroids.
@@
where H is the representation of all samples and $\textbf{h}_{i}$ is the representation of the $i$-th sample, $\mathcal{C}$ denotes the cluster centroid set, and $\textbf{c}_{k}\in\mathcal{C}$ represents the $k$-th cluster centroid. The sample representation associated with the $k$-th cluster centroid is expressed as $\textbf{h}_{j}\in\textbf{H}_{\textbf{c}_{k}}$, while $\textbf{H}_{\textbf{c}_{k}}$ denotes the sample representations within the $k$-th cluster centroid.
VendiScore. VendiScore (Dan Friedman & Dieng, 2023) is a recently proposed diversity metric that falls under the category of the transformation-based method. Based on sample representations, VendiScore utilizes a kernel function to calculate similarity matrix K. Subsequently, VendiScore summarizes diversity as the exponential of the Shannon entropy of the eigenvalues of $\textbf{K}/n$.
@@
where $\operatorname{Kernel}(\cdot)$ is the kernel function, such as the cosine similarity, $\lambda_{i}$ is the $i$-th eigenvalue of $\textbf{K}/n$.

### v2

@@ C.2.1 Hyperparameter Settings of Diversity Evaluation
  • DCScore. We employ the inner product as $\operatorname{Kernel}(\cdot)$, and Softmax as $f_{\textbf{K}}(\cdot)$. Except for hyperparameter sensitivity experiments, we set $\tau$ in Eq. (4) to 1 for all other experiments.
@@
  • VendiScore. We employ the inner product as $\operatorname{Kernel}(\cdot)$.
C.2.2 Hyperparameter Settings of Downstream Task Training
@@ C.2.3 Evaluation protocol
C.2.4 Prompt Settings of LLM Evaluation
In Section 5.2.3, we use GPT-4 to perform pairwise diversity comparisons. To guide GPT-4 in making these comparisons, we employ a well-designed prompt, as illustrated in Figure 7. The prompt for GPT-4 evaluations includes the task definition, diversity definition, general prompt, and sentence sets to be compared.
@@ C.2.4 Prompt Settings of LLM Evaluation
C.2.5 Settings of Human Evaluation
According to Appendix C.1.1, datasets in Section 5.2.2 are generated at six temperatures, with five results per context (prompt) in each sub-dataset. During the evaluation, evaluators are asked to select the more diverse sub-dataset from pairs of sub-datasets. Across six temperatures, this results in 15 comparisons, and with three evaluators, a total of 45 judgments are made. Sub-datasets are ranked by the frequency of being chosen as more diverse. This process is repeated five times with different contexts to derive the final human diversity ranking.
Appendix D Additional Experiments
D.1 Computational Cost for Larger Datasets
In the case where the inner product is used as the kernel function and $n\gg d$, VendiScore can significantly reduce computational complexity. To ensure a fair comparison, we compare the computation times of VendiScore and DCScore on larger datasets, as well as under different kernel functions. Specifically, we employ SST2, Yelp, and AG News as the evaluated datasets. We randomly select 4k, 8k, 16k, 32k, 64k samples and record the computation times for both methods across these different sample sizes. We repeat the experiments 5 times to report the mean and standard deviation.
As shown in Tables 610, and 11, DCScore has a shorter computation time than VendiScore in most cases, with VendiScore only exhibiting a computational advantage when the inner product is used and $n\gg d$. Furthermore, as the sample size increases, the efficiency advantage of DCScore becomes more pronounced. When using a polynomial kernel, on the SST2 dataset, DCScore requires only one-third of the computation time of VendiScore when the sample size reaches 64k. In contrast, although VendiScore has a computational advantage in the case of the inner product, the difference compared to DCScore is not significant. The experimental results are consistent with our complexity analysis presented in Section 4.3. Overall, DCScore outperforms VendiScore in terms of computation time across most kernel functions. VendiScore exhibits a computational time advantage only when the inner product is used as the kernel function, which will limit its applicability. As shown in Chapter 4 of (Seeger, 2004), it is essential to employ different kernel functions to accommodate a wider range of scenarios.
@@ D.1 Computational Cost for Larger Datasets
D.2 Comparison of Computational Cost between DCScore and Efficiency-improved Methods
A recent work (Ospanov et al., 2024) employs the random Fourier features framework to reduce the computational cost of VendiScore (Dan Friedman & Dieng, 2023) and RKE (Jalali et al., 2023). To further investigate the computational efficiency of DCScore, we compare it with efficiency-improved versions of two entropy-based diversity evaluation methods, namely FKEA-Vendi and FKEA-RKE. Here, we set $\alpha$ to 1 for FKEA-Vendi and 2 for FKEA-RKE. Notably, we follow the experimental settings described in Appendix D.1 and leverage different random seeds for data sampling. Table 12 illustrates the computation time comparison across the SST2, Yelp, and AG News datasets. DCScore demonstrates lower computation time compared to FKEA-Vendi and FKEA-RKE on the SST2 and AG News datasets. However, the results are reversed for the Yelp dataset. Additionally, we observe that FKEA-Vendi does not improve computation time compared to VendiScore, possibly because FKEA focuses solely on reducing the complexity of the summarization (eigenvalue computation) phase, whereas our computation time encompasses the entire calculation process.
Table 12: Comparison of computation time between DCScore and two efficiency-improved methods (FKEA-Vendi and FKEA-RKE).
Datasets Sample num 4k 8k 16k 32k 64k
SST2 FKEA-Vendi 18.43±0.14 36.97±0.03 74.03±0.20 147.78±0.30 295.14±0.41
FKEA-RKE 18.46±0.08 37.08±0.12 73.98±0.09 147.99±0.09 297.46±2.02
DCScore 3.26±0.08 6.75±0.08 13.94±0.06 31.32±0.34 90.36±1.24
Yelp FKEA-Vendi 26.28±0.13 52.69±0.24 106.59±2.25 212.52±2.22 422.51±1.19
FKEA-RKE 26.19±0.13 52.68±0.24 105.27±0.22 212.38±1.87 422.94±1.55
DCScore 37.53±0.08 75.19±0.10 151.59±0.39 306.80±0.33 641.60±0.60
AG News FKEA-Vendi 20.64±0.05 41.66±0.06 83.02±0.11 165.92±0.25 338.16±14.26
FKEA-RKE 20.68±0.04 41.68±0.03 83.04±0.14 165.85±0.23 331.58±0.29
DCScore 10.29±0.49 21.12±0.73 43.84±1.02 91.64±0.86 213.30±2.15
D.3 Correlation with Downstream Task Training
To investigate the correlation between DCScore and downstream task training, we train text classification models using self-generated datasets under zero-shot and few-shot settings. We vary the generation temperature $\tau_{g}$ of self-generated datasets within the range of $\{0.2,0.7,1.2\}$. More details of training datasets and hyperparameters are presented in Appendix C.1.1 and C.2.2, respectively. Figure 8 shows the loss curves of these trained classification models. In the zero-shot setting, we observe increasing optimal loss values as $\tau_{g}$ varied from 0.2 to 1.2, indicating that the model is more easily fitted to datasets with limited diversity. However, as shown in Table 7, models trained on more diverse datasets achieve better accuracy, which can be attributed to their enhanced generalization capabilities. From Table 7, the diversity evaluated by DCScore has a similar trend to the accuracy performance in the zero-shot setting, further demonstrating the effectiveness of DCScore in diversity evaluation.
@@ D.3 Correlation with Downstream Task Training
In the few-shot setting, we observe a trend in optimal loss variation similar to that in the zero-shot setting, as shown in Figure 8. However, Figure 9 reveals that at epoch 120, models trained on datasets generated with $\tau_{g}=0.7$ outperform those using $\tau_{g}=1.2$. This phenomenon can be attributed to the higher diversity of datasets generated at a higher $\tau_{g}$, resulting in increased fitting difficulty. Under the current settings, the number of training epochs for the dataset generated at a temperature of 1.2 is insufficient, preventing the trained model from achieving optimal performance. To validate this hypothesis, we increase the number of epochs to 240 and 360 and train models on the dataset generated at a temperature of 1.2. The final training loss and accuracy of these models are shown in Figure 9. We observe that as the number of epochs increases, the model’s loss gradually decreases, and its performance improves progressively. Ultimately, the model’s accuracy outperforms that of models trained on datasets generated at temperatures of 0.2 and 0.7. Moreover, from Table 7, models trained on datasets from the zero-shot setting outperform those trained on datasets from the few-shot setting. However, this discrepancy arises from the different test sets used in the two settings, making direct performance comparisons inappropriate.
@@
D.4 Impact of Embedding Functions $\Phi$
The paradigm of the transformation-based method enables DCScore to utilize various embedding functions tailored to different scenarios. Consequently, we investigate the impact of embedding functions on self-generated datasets used in Section 5.2.1. As shown in Table 13, we compare the correlation of the diversity evaluation results of DCScore across 4 different embedding functions with diversity pseudo-truths, where the model names in parentheses within the embedding function refer to those available on Hugging Face. Our findings indicate that DCScore exhibits strong correlations with diversity pseudo-truths across various embedding functions. Notably, DCScore utilizing the BGE embedding function achieves the best results in half of the cases. Additionally, the minimum correlation in Table 13 exceeds 0.96, which is classified as a strong correlation according to (Akoglu, 2018). This result also supports the following two conclusions: (1) the embedding function used effectively captures the differences among samples from multiple perspectives, and (2) DCScore is sufficiently adaptable to different embedding functions while maintaining stable performance.
Table 13: Correlation (Spearman’s $\rho$) results of DCScore with various embedding functions. Spearman’s $\rho$ varies between -1 and +1, with 0 implying no correlation. Best results are indicated in bold.
@@ D.4 Impact of Embedding Functions $\Phi$
D.5 Impact of Kernel Functions
Similar to Appendix D.4, we investigate the impact of different kernel functions on the performance of DCScore. Specifically, this experimental setup is identical to that in Appendix D.4. As shown in Table 14, we find that DCScore demonstrates stable performance across various kernel functions. However, the influence of the kernel function is slightly more pronounced than that of the embedding function, as indicated by the greater fluctuations in correlation among the different kernel functions. Furthermore, we observe that DCScore achieves optimal performance in the case of the inner product. Overall, DCScore consistently maintains strong diversity evaluation performance across different kernel functions.
Table 14: Correlation (Spearman’s $\rho$) results of DCScore with various kernel functions. Spearman’s $\rho$ varies between -1 and +1, with 0 implying no correlation. Best results are indicated in bold.
@@ D.5 Impact of Kernel Functions
Appendix E Baseline Methods
We present the detailed modeling of baseline methods into DCScore as follows:
Distinct-n. Distinct-n (Li et al., 2015) is a prevalent diversity metric depending on n-grams, where n signifies the number of successive items. Distinct-n calculates the proportion of unique n-grams to all n-grams. The n-grams operation falls under the text representation stage, while the step of obtaining a unique set of n-grams corresponds to the pairwise similarity stage. Typically, a high form similarity among samples in the evaluated dataset results in a smaller unique n-gram set. Finally, ratio calculations belong to the diversity summarization stage.
@@ Appendix E Baseline Methods
K-means inertia. K-means Inertia (Du & Black, 2019), a transformation-based method, performs clustering in sample representation and then calculates inertia as diversity outcomes. Here, inertia refers to the square summation of the distance between samples and cluster centroids.
@@
where H is the representation of all samples and $\textbf{h}_{i}$ is the representation of the $i$-th sample, $\mathcal{C}$ denotes the cluster centroid set, and $\textbf{c}_{k}\in\mathcal{C}$ represents the $k$-th cluster centroid. The sample representation associated with the $k$-th cluster centroid is expressed as $\textbf{h}_{j}\in\textbf{H}_{\textbf{c}_{k}}$, while $\textbf{H}_{\textbf{c}_{k}}$ denotes the sample representations corresponding to the $k$-th cluster centroid.
VendiScore. VendiScore (Dan Friedman & Dieng, 2023) is a recently proposed diversity metric that falls under the category of the transformation-based method. Based on sample representations, VendiScore utilizes a kernel function to calculate a similarity matrix K. Subsequently, VendiScore summarizes diversity as the exponential of the Shannon entropy of the eigenvalues of $\textbf{K}/n$.
@@
where $\operatorname{Kernel}(\cdot)$ is the kernel function, such as the inner product, $\lambda_{i}$ is the $i$-th eigenvalue of $\textbf{K}/n$.